When Y was born, a bank account was started in Y’s name with $20.05 The account has earned an average annual rate of return of 4%. Today, the account is valued at $500.00. How old is Y?

1
Expert's answer
2021-09-28T18:10:40-0400

Calculation of the age of Y:

The age of Y will be equal to the time it will take $20.05 to grow to $500 at the average annual rate of return of 4%:

Formula,

"Future value = Present value x (1+ rate)^n"

"500 = 20.05 x (1+ 0.04)^n"

"500 \/ 20.05 = (1.04)^n"

"24.937656 = (1.04)^n"

Apply log both the side,

"Log(24.937656) = Log(1.04)^n"

"Log(24.937656) = nLog(1.04)"

"n = Log(24.937656) \/ Log(1.04)"

"n = 1.396856 \/ 0.017033"

"n = 82"

So, it will take 82 years to $20.05 to grow to $500 at the average annual rate of return of 4%. Hence, Y is 82 years old now.